OpenStudy (anonymous):

What is the geometric mean of 3 and 15? 45 5 radical 3 3 radical 5 22.5

OpenStudy (anonymous):

the geometric mean of \(a\) and \(b\) is \(\sqrt{ab}\) in your example a first step is \(\sqrt{3\times 15}\)

OpenStudy (anonymous):

so wait sqrt means? square ...or square root?

OpenStudy (anonymous):

square is \(x^2\) square root is \(\sqrt{x}\)

OpenStudy (anonymous):

if you are looking literally at "sqrt" refresh your browser to see the symbols you should see symbols, not code

OpenStudy (anonymous):

I still see the same even after I refreshed the browser. okay so im suppose to squareroot is 3 times it 15 and I will get my answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get 25.9 as my answer and if i round it, its 29.10 but that's not a answer choice